# A Survey on Cross-Domain Sequential Recommendation

### v3

@@ 1 Introduction
In real life, people leave interaction traces in multiple scenarios (e.g., shopping scenarios, reading scenarios, etc.), multiple platforms (e.g., Amazon, Taobao, etc.), and even multiple boards (e.g., books and movies in Douban, etc.) all the time. Considering all these as different domains, it is a critical step for developing recommender systems to combine information across domains so that the data of user preferences is no longer fragmented, shifting the modeling from flat to stereoscopic. As shown in Table 1, recommender systems focus on information that is becoming deeper and more comprehensive, starting from traditional one-class collaborative filtering (OCCF), further progressing to sequential one-class collaborative filtering (SOCCF) and cross-domain one-class collaborative filtering (CD-OCCF), and finally to cross-domain sequential one-class collaborative filtering (CD-SOCCF, also called cross-domain sequential recommendation, CDSR).
@@
In the case of OCCF Rendle et al. (2012), the available information is limited to whether a user interacts with an item. As illustrated in Figure 1, we only know all the items that user 1 and user 2 have interacted with in a single domain, without considering the sequential pattern among successive items. SOCCF Lin et al. (2020), building upon OCCF, places greater emphasis on the order relationships between items within a sequence, i.e., the information of intra-sequence in a single domain, and thus the model can capture the user’s long-term and short-term preferences. At a large granularity, expanding from a single domain to multiple domains offers a fresh perspective to address the data sparsity issue in recommender systems, thus constituting CD-OCCF Hu et al. (2018). For instance, as shown in Figure 1, users have interactions not only in the book domain but also in the closely related movie domain. In a cross-domain scenario, we must meet the challenges of transferring or aggregating interaction records from multiple domains and aligning representations of users or items across different domains.
Indeed, CDSR combines all of the aforementioned directions and further incorporates sequential information based on CD-OCCF and cross-domain information based on SOCCF. As indicated by the purple arrows in Figure 1, CDSR necessitates the precise capturing of sequential relationships between items across different domains. For example, user 1, after reading two books in the $Harry\ Potter$ series, chooses to watch the corresponding $Harry\ Potter$ movie. Similarly, user 2, who enjoys watching romantic movies, also reads the original novels after watching the $Pride\ and\ Prejudice$ movie. Besides, it is also worth considering in CDSR how to fuse sequential information from different domains as well as how to distinguish between users’ specific preferences within a single domain and global preferences shared across multiple domains.
In this survey, we first formulate the CDSR problem and modeling tasks, considering various dimensionality reductions and different input representations. Then we adopt both macro and micro views to summarize the existing works in CDSR. From a macro view, we present an overview of multi-level fusion structures, discussing how to fuse information across different domains and exploring bridges for cross-domain fusion. From a micro view, we conduct a detailed analysis of various technologies employed by existing works that are categorized into basic and auxiliary learning technologies. Based on this analysis, we obtain a comprehensive classification shown in Table 2 and framework figure of the key technologies ( shown in Figure 5). Furthermore, we list the datasets commonly used in CDSR and the representative experimental results as well as provide some insights into potential future directions.
@@ 2.1 Problem Definition
As temporal information and cross-domain information are incorporated, the interaction data between users and items in CDSR undergoes a gradual expansion into a four-dimensional data tensor, consisting of dimensions about users, items, time, and domains. We denote this four-dimensional tensor with $\Gamma\in\mathbb{R}^{n\times m\times s\times k}$, where $n$ is the number of users, $m$ is the number of items, $s$ is the discrete time intervals, and $k$ is the number of domains. Each element $\gamma_{(u,i,t,d)}$ in it signifies whether there is an interaction between user $u$ and item $i$ at time $t$ in domain $d$. The goal of CDSR is to estimate the probability for all candidate items in each domain, and then to recommend the most probable next item to each user. The estimated probability can be formalized as follows,
@@
where $\hat{i}$ denotes the candidate item in each domain, $\Gamma$ is the four-dimensional tensor containing all interaction information, and $f(\ \Gamma\ )$ indicates the learned function to estimate $P(\ \hat{i}\ |\ \Gamma\ )$.
@@ 2.2 Multidirectional Dimensionality Reduction
For computational and modeling convenience, the four-dimensional tensor is often reduced to some three-dimensional tensors from various directions ( shown in Figure 2). Below, we explain dimensionality reduction from three directions:
@@
Indeed, these dimensionality reduction directions are often combined and used in conjunction rather than being treated as independent. Next, we visualize these reduction processes on different input representations.
@@ 2.3.3 Side Information
More and more researchers consider incorporating side information to enrich the semantic representation of users’ historical behaviors. Here, we divide the most common side information into three categories: time, text (e.g., user/item context and review), and knowledge graph.
Time. Previously the timestamps are only used to order the items, but some researchers Xiao et al. (2023); Wang et al. (2022); Guo et al. (2022) define $t_{ij}=|t_{i}-t_{j}|$ to model the time interval and apply it to the subsequent learning of a model, where $t_{i}$ and $t_{j}$ are the timestamps that the item $i$ and the item $j$ are interacted with, respectively. Besides, we can also mine more information about time, such as periodicity and the duration of the interaction, etc.
User/Item Context and Review. Except for the user ID, triples composed of multiple contextual information (e.g., (``$UserID",``City",``Age",\cdots)$) are utilized as a basis for finding user-user relationships Ouyang et al. (2020). Similarly, the context information that consists of categories, tags, keywords, etc, is used as a supplement to item ID Xiao et al. (2023); Ouyang et al. (2020); Zhuang et al. (2020). Review as a type of available text message can associate users and items well. A common way to encode these texts is through some pre-trained models.
Knowledge Graph. Knowledge graphs Bi et al. (2020); Ma et al. (2022) deal with information from both entity and relationship views. It not only contains structural information between nodes but also implies some relationship descriptions. A knowledge graph defines an entity set $E^{KG}$ and a relation set $R^{KG}$, which consists of multiple entity-relation-entity triples $<e_{i},r,e_{j}>$ (e.g., $<e_{i_{1}^{A}},Is\_the\_same\_category,e_{i_{1}^{B}}>$ meaning that the entity $e_{i_{1}^{A}}$ from domain A has the same category as entity $e_{i_{1}^{B}}$ from domain B).
@@ 3 Macro-View: What Structure Is Used to Fuse the CDSR Information?
In CDSR, the fusion structure serves as the primary skeleton of a model, providing pathways for separating and aggregating these features. In this section, we first describe multi-level fusion structures and then elaborate on the bridge for inter-domain fusion.
@@ 3.1.1 One-Level Fusion Structures
In earlier works, to combine the cross-domain information, there are a lot of works that first learn user preferences from domain A and domain B separately, and then fuse the representations of the two domains using various operations (shown in Figure 3(a)). These operations include, but are not limited to, concatenation Lei et al. (2021); Guo et al. (2023c), summation Alharbi and Caragea (2021); Ding et al. (2023), multi-layer perceptron (MLP) Bi et al. (2020), and some attention mechanisms Ouyang et al. (2020); Li et al. (2021).
Additionally, some works employ transfer learning to transfer knowledge from a source domain to a target domain Chen et al. (2021); Liu and Zhu (2021), or train a discriminator to bridge the representations of two domains with the idea of adversarial learning Li et al. (2022), shown in Figure 3(b).
@@ 3.1.2 Two/Three-Level Fusion Structures
Considering the potential for further exploration and advancement of fusion structures, some works combine the single domain and the hybrid domain with a multi-level fusion structure. As shown in Figure 3(e), in order to predict the next item in a separate domain, the hybrid domain is utilized as the main sharer Cao et al. (2022); Wang et al. (2022) or the bridge Zheng et al. (2022) to transfer knowledge from another domain. There are also some works Ye et al. (2023) that choose to combine domain A and domain B first and then fuse the hybrid information (i.e., Figure 3(f)).
@@ 3.1.3 Discussion
Although there are various multi-level fusion structures, it does not mean that a more complex structure will result in a better performance. The simple fusion structures, i.e., one-level fusion structures, are easy to implement but may not be conducive to comprehensive modeling of domain-specific and domain-generic features due to the limitation of the degree of fusion. While improving effectiveness, multi-level fusion structures bring increased complexity and reduced interpretability.
Focusing on the symmetry and asymmetry of the structures, the design of a cross-domain fusion structure also has a lot to do with the extent to which the researchers focus on each domain. For example, some researchers Ouyang et al. (2020); Bi et al. (2020); Xu et al. (2023c) consider the two domains as source and target domains and leverage the data-rich source domain to assist the data-sparse target domain, which makes the target domain dominant in the fusion structure. Other researchers Guo et al. (2021); Li et al. (2021); Ma et al. (2022) aim to improve the recommendation performance of both domains at the same time, which makes the fusion structure designed tend to be more symmetric.
@@ 3.2 Bridges for Cross-Domain Fusion
  • The Same Users. If some same users exist, then they are often the first choice as the bridge. Some works only focus on the fully overlapping users, neglecting non-overlapping users, while others Li et al. (2022) consider both cases.
@@
  • Clusters/Groups. Further restricting the condition to no overlapping users and no side information, we can work from the clusters/groups, since a specific group of users could have similar preferences Lin et al. (2023a).
Different bridges exhibit distinct characteristics, each suitable for different scenarios. Works relying on overlapping users often perform well but their application scenarios may face limitations due to the sparse real-world data. Conversely, works built on non-overlapping users can be applied to a broader range of scenarios but may have limited performance.
Table 2: A systematic overview of the existing models for CDSR.
Data Structure Basic Technology Auxiliary Learning Paper
GNN Attention RNN Contrastive Learning Transfer Learning Other
Sequence DREAM Ye et al. (2023)
P-CDSR Xiao et al. (2023)
Tri-CDR Ma et al. (2023)
SEMI Lei et al. (2021)
CGRec Park et al. (2023)
LCN Hou et al. (2023)
MACD Xu et al. (2023a)
Adversarial Learning TPUF Ding et al. (2023)
RecGURU Li et al. (2022)
DA-DAN Guo et al. (2023a)
Reinforcement Learning RL-ISN Guo et al. (2023c)
Prompt Learning PLCR Guo et al. (2023b)
SATLR Liu and Zhu (2021)
MAN Lin et al. (2023a)
MiNet Ouyang et al. (2020)
CD-ASR Alharbi and Caragea (2021)
CD-SASRec Alharbi and Caragea (2022)
CMVCDR Zang et al. (2023)
DASL Li et al. (2021)
CDNST Zhuang et al. (2020)
PSJNet Sun et al. (2023)
$\pi$-Net Ma et al. (2019)
CDHRM Wang et al. (2020)
SCLSTM Yang et al. (2020)
MSECDR Hong and Jung (2023)
AMID Xu et al. (2023b)
Graph DA-GCN Guo et al. (2021)
TiDA-GCN Guo et al. (2022)
EA-GCL Wang et al. (2023)
MGCL Xu et al. (2023c)
C2DSR Cao et al. (2022)
Federated Learning FedDCSR Zhang et al. (2023a)
DDGHM Zheng et al. (2022)
LEA-GCN Zhang et al. (2023b)
IESRec Liu et al. (2023)
DAT-MDI Chen et al. (2021)
SGCross Li et al. (2023)
AGNNGRU-CDR Qu et al. (2021)
DCDIR Bi et al. (2020)
MIFN Ma et al. (2022)
CsrGCF Wang et al. (2022)
@@ 4 Micro-View: What Technologies Are Used to Address the CDSR Problem?
In this section, we take a more tangible perspective to summarize the technologies adopted by the existing models in addressing the challenges. In conjunction with Table 2 and Figure 5, we elaborate on the basic technologies and the auxiliary learning technologies, respectively.
@@ 4.1.2 Attention Mechanisms in CDSR
Attention mechanisms can selectively focus on the importance of different parts and weight their contributions accordingly, enabling effective aggregation. Attention mechanisms are utilized in two main ways in CDSR modeling. One is to use an attention-based encoder (e.g., Transformer Vaswani et al. (2017), SASRec Kang and McAuley (2018) or multi-head attention blocks, etc.), replacing RNNs as a sequence encoder Ye et al. (2023); Xu et al. (2023c); Alharbi and Caragea (2022); Ding et al. (2023); Park et al. (2023). The other is to obtain learnable attention weights thus aggregating cross-domain information at multiple levels. For instance, MiNet Ouyang et al. (2020) designs item-level attention and interest-level attention to learn which items are more important and which of these items better matches a certain interest of the user. With representation-level attention, C2DSR Cao et al. (2022) fuses the sequential representation and the graph representation together. By adopting domain-level attention, Tri-CDR Ma et al. (2023) learns the weights of the features from different domains and then aggregates them. And MAN Lin et al. (2023a) also proposes a group-level attention to bridge the not-aligned information. In fact, attention strategies need to be designed from multiple levels which poses a challenge to researchers. However, it needs to use the interaction of all the positions when calculating the weights, so it contains a large number of parameters and may suffer from the issue of data sparsity.

### v4

@@ 1 Introduction
In real life, people always leave interaction traces in multiple scenarios (e.g., shopping scenarios, reading scenarios, etc.), multiple platforms (e.g., Amazon, Taobao, etc.), and even multiple boards (e.g., books and movies in Douban, etc.). Considering all these as different domains, combining in-depth information across domains is a critical step for the development of recommender systems. It makes the data of user preferences no longer fragmented and shifts the modeling from flat to stereoscopic. As shown in Table 1, recommender systems focus on information that is becoming deeper and more comprehensive, starting from traditional one-class collaborative filtering (OCCF), further progressing to sequential one-class collaborative filtering (SOCCF) and to cross-domain one-class collaborative filtering (CD-OCCF), and finally to cross-domain sequential one-class collaborative filtering (CD-SOCCF, also called cross-domain sequential recommendation, CDSR).
@@
In the case of OCCF Rendle et al. (2012), the available information is limited to whether a user interacts with an item. As illustrated in Figure 1, we only know all the items that User 1 and User 2 have interacted with in a single domain, without considering the sequential pattern among successive items. SOCCF Lin et al. (2020), building upon OCCF, places greater emphasis on the order relationships between items within a sequence, i.e., the information of intra-sequence in a single domain, which enables the model to capture the user’s long-term and short-term preferences. At a large granularity, expanding from a single domain to multiple domains offers a fresh perspective to address the data sparsity issue in recommender systems, thus constituting CD-OCCF Hu et al. (2018). For instance, as shown in Figure 1, users have interactions not only in the book domain but also in the closely related movie domain. In a cross-domain scenario, we must meet the following challenges: how to transfer or aggregate interaction records from multiple domains and how to align representations of users or items across different domains Zhu et al. (2021b).
Indeed, CDSR combines all of the aforementioned directions and further incorporates sequential information based on CD-OCCF and cross-domain information based on SOCCF. As indicated by the purple arrows in Figure 1, CDSR necessitates the precise capturing of sequential relationships between items across different domains. For example, User 1, after reading two books in the $Harry\ Potter$ series, chooses to watch the corresponding $Harry\ Potter$ movie. Similarly, User 2 who enjoys watching romantic movies also reads the original novels after watching the $Pride\ and\ Prejudice$ movie. Besides, it is also worth considering in CDSR how to fuse sequential information from different domains as well as how to distinguish between users’ specific preferences within a single domain and global preferences shared across multiple domains.
In this survey, we first formulate the CDSR problem and modeling tasks, considering various dimensionality reductions and different input representations. Then we adopt both macro and micro views to summarize the existing works in CDSR. From a macro view, we present an overview of multi-level fusion structures, discussing how to fuse information across different domains and exploring bridges for cross-domain fusion. From a micro view, we conduct a detailed analysis of various technologies employed by existing works that are categorized into basic and auxiliary learning technologies. Based on this analysis, we obtain a comprehensive classification (shown in Table 2) and framework figure of the key technologies (shown in Figure 5). Furthermore, we list the datasets commonly used in CDSR and the representative experimental results as well as provide some insights into potential future directions.
@@ 2.1 Problem Definition
As temporal information and cross-domain information are incorporated, the interaction data between users and items in CDSR undergoes a gradual expansion into a four-dimensional data tensor, consisting of dimensions about users, items, time, and domains. We denote this four-dimensional tensor with $\Gamma\in\mathbb{R}^{n\times m\times s\times k}$, where $n$ and $m$ is the number of users and items in all domain, $s$ is the discrete time intervals, and $k$ is the number of domains. Each element $\gamma_{(u,i,t,d)}$ in it signifies whether there is an interaction between user $u$ and item $i$ at time $t$ in domain $d$. The goal of CDSR is to estimate the probability for all candidate items in each domain, and then to recommend the most probable next item to each user. The estimated probability can be formalized as follows,
@@
where $\hat{i}$ denotes the candidate item of each user in each domain, $\Gamma$ is the four-dimensional tensor containing all interaction information, and $f(\ \Gamma\ )$ indicates the learned function to estimate $P(\ \hat{i}\ |\ \Gamma\ )$.
@@ 2.2 Multidirectional Dimensionality Reduction
For computational and modeling convenience, the four-dimensional tensor is often reduced to some three-dimensional tensors from various directions (as shown in Figure 2). Below, we explain dimensionality reduction from three directions:
@@
Indeed, these dimensionality reduction directions are often combined and used in conjunction rather than being treated as independent. Adopting only one reduction direction may result in missing information to a certain degree. Next, we visualize these reduction processes on different input representations.
@@ 2.3.3 Side Information
More and more researchers consider incorporating side information to enrich the semantic representation of users’ historical behaviors. In this section, we divide the most common side information into three categories: time, text (e.g., user/item contexts and reviews), and knowledge graph.
Time. Previously the timestamps are only used to order the items, but some researchers Xiao et al. (2023); Wang et al. (2022); Guo et al. (2022) define $t_{ij}=|t_{i}-t_{j}|$ to model the time interval and apply it to the subsequent learning of model, where $t_{i}$ and $t_{j}$ are the timestamps that the item $i$ and the item $j$ are interacted with, respectively. Besides, we can also mine more information about time, such as periodicity and the duration of the interaction, etc.
User/Item Contexts and Reviews. Except for the user ID, triples composed of multiple contextual information (e.g., (``$UserID",``City",``Age",\cdots)$) are utilized as a basis for finding user-user relationships Ouyang et al. (2020). Similarly, the context information that consists of categories, tags, keywords, etc, is used as a supplement to item ID Xiao et al. (2023); Ouyang et al. (2020); Zhuang et al. (2020). Review as a type of available text message can associate users and items well. A common way to encode these texts is through some pre-trained models, such as BERT Devlin et al. (2018).
Knowledge Graph. Knowledge graphs Bi et al. (2020); Ma et al. (2022) deal with information from both entity and relationship views. It not only contains structural information between nodes but also implies some relationship description. A knowledge graph defines an entity set $E^{KG}$ and a relation set $R^{KG}$, which consists of multiple entity-relation-entity triples $<e_{i},r,e_{j}>$ (e.g., $<e_{i_{1}^{A}},Is\_the\_same\_category,e_{i_{1}^{B}}>$ meaning that the entity $e_{i_{1}^{A}}$ from domain A has the same category as entity $e_{i_{1}^{B}}$ from domain B).
@@ 3 Macro-View: What Structure Is Used to Fuse the CDSR Information?
In CDSR, the fusion structure serves as the primary skeleton of a model, providing pathways for the separation and aggregation of these features. In this section, we first describe multi-level fusion structures and then elaborate on the bridge for inter-domain fusion.
@@ 3.1.1 One-Level Fusion Structures
In earlier works, to combine the cross-domain information, there are a lot of works that first learn user preferences from domain A and domain B separately, and then fuse the representations of the two domains using various operations (as shown in Figure 3(a)). These operations include, but are not limited to, concatenation Lei et al. (2021); Guo et al. (2023c), summation Alharbi and Caragea (2021); Ding et al. (2023), multi-layer perceptron (MLP) Bi et al. (2020), and some attention mechanisms Ouyang et al. (2020); Li et al. (2021).
Additionally, some works employ transfer learning to transfer knowledge from a source domain to a target domain Chen et al. (2021); Liu and Zhu (2021), or train a discriminator to bridge the representations of two domains with the idea of adversarial learning Li et al. (2022), as shown in Figure 3(b).
@@ 3.1.2 Two/Three-Level Fusion Structures
Considering the potential for further exploration and advancement in fusion structures, some works combine the single domain and the hybrid domain with a multi-level fusion structure. As shown in Figure 3(e), in order to predict the next item in a separate domain, the hybrid domain is utilized as the main sharer Cao et al. (2022); Wang et al. (2022) or the bridge Zheng et al. (2022) to transfer knowledge from another domain. There are also some works Ye et al. (2023) that choose to combine the domain A and domain B first and then fuse the hybrid information (i.e., Figure 3(f)).
@@ 3.1.3 Discussion
Although there are various multi-level fusion structures, it does not mean that a more complex structure will perform better. Limited by the degree of fusion, the simple fusion structures, i.e., one-level fusion structures, are easy to implement but may not comprehensively model the domain-specific and domain-generic features. While improving effectiveness, multi-level fusion structures bring increased complexity and reduced interpretability.
Considering the symmetry and asymmetry of the structures, which the researchers focus on each domain greatly affects the design of a cross-domain fusion structure. For example, some researchers Ouyang et al. (2020); Bi et al. (2020); Xu et al. (2023c) consider the two domains as source and target domains and leverage the data-rich source domain to assist the data-sparse target domain, which makes the target domain dominant in the fusion structure. Other researchers Guo et al. (2021); Li et al. (2021); Ma et al. (2022) aim to improve the recommendation performance of both domains simultaneously, making the fusion structure designed tend to be more symmetric.
@@
Table 2: A systematic overview of the existing models for CDSR.
Data Structure Basic Technology Auxiliary Learning Paper
GNN RNN Attention
Contrastive
Learning
Transfer
Learning
Others
Sequence
$\pi$-Net Ma et al. (2019), CDHRM Wang et al. (2020),
SCLSTM Yang et al. (2020), PSJNet Sun et al. (2023)
CDNST Zhuang et al. (2020)
DASL Li et al. (2021), TJAPL Xu et al. (2024)
CMVCDR Zang et al. (2023)
SEMI Lei et al. (2021), DREAM Ye et al. (2023), P-CDSR Xiao et al. (2023),
Tri-CDR Ma et al. (2023), CGRec Park et al. (2023), LCN Hou et al. (2023),
MACD Xu et al. (2023a)
MiNet Ouyang et al. (2020), CD-ASR Alharbi and Caragea (2021),
CD-SASRec Alharbi and Caragea (2022), MAN Lin et al. (2023a)
SATLR Liu and Zhu (2021)
Adversarial
Learning
RecGURU Li et al. (2022), TPUF Ding et al. (2023), DA-DAN Guo et al. (2023a)
Reinforcement
Learning
RL-ISN Guo et al. (2023c), O-SCDR Nanthini and Kumar (2024)
Prompt
Learning
PLCR Guo et al. (2023b)
MSECDR Hong and Jung (2023), AMID Xu et al. (2023b)
Graph DCDIR Bi et al. (2020), MIFN Ma et al. (2022)
AGNNGRU-CDR Qu et al. (2021)
DAT-MDI Chen et al. (2021), SGCross Li et al. (2023)
DA-GCN Guo et al. (2021), TiDA-GCN Guo et al. (2022)
C2DSR Cao et al. (2022), EA-GCL Wang et al. (2023), MGCL Xu et al. (2023c)
Federated
Learning
FedDCSR Zhang et al. (2023a)
DDGHM Zheng et al. (2022)
LEA-GCN Zhang et al. (2023b), IESRec Liu et al. (2023)
CsrGCF Wang et al. (2022)
@@ 3.2 Bridges for Cross-Domain Fusion
  • Same Users. If some same users exist, then they are often the first choice as the bridge. Some works only focus on the fully overlapping users, neglecting non-overlapping users, while others Li et al. (2022) consider both cases.
@@
  • Clusters/Groups. Further restricting the condition to no overlapping users and no side information, the clusters/groups are another entry point since a specific group of users could have similar preferences Lin et al. (2023a).
Works relying on overlapping users often perform well but their application scenarios may face limitations due to the sparse real-world data. Conversely, works built on non-overlapping users can be applied to a broader range of scenarios but may have limited performance. In fact, cross-domain can be built on more than one type of bridge.
@@ 4 Micro-View: What Technologies Are Used to Address the CDSR Problem?
In this section, we take a more tangible perspective to summarize the technologies adopted by the existing models in addressing the challenges in CDSR. In conjunction with Table 2 and Figure 5, we elaborate on the basic technologies and the auxiliary learning technologies, respectively. In Figure 6, we give some examples of the utilization of the base technologies.
@@ 4.1.2 Attention Mechanisms in CDSR
Attention mechanisms can selectively focus on the importance of different parts and weight their contributions accordingly, enabling effective aggregation. Attention mechanisms are utilized in two main ways in CDSR. One is to use an attention-based encoder (e.g., Transformer Vaswani et al. (2017), SASRec Kang and McAuley (2018) or multi-head attention blocks, etc.), replacing RNNs as a sequence encoder Ye et al. (2023); Xu et al. (2023c); Alharbi and Caragea (2022); Ding et al. (2023); Park et al. (2023). The other is to obtain learnable attention weights thus aggregating cross-domain information at multiple levels. For instance, MiNet Ouyang et al. (2020) designs item-level attention and interest-level attention to learn which items are more important and which of these items better matches a certain interest of the user. With representation-level attention, C2DSR Cao et al. (2022) fuses the sequential representation and the graph representation together. By adopting domain-level attention, Tri-CDR Ma et al. (2023) learns the weights of the features from different domains and then aggregates them. And MAN Lin et al. (2023a) also proposes a group-level attention to bridge the not-aligned information. In fact, attention strategies need to be designed from multiple levels which poses a challenge to researchers. However, it needs to use the interaction of all the positions when calculating the weights, so it contains a large number of parameters and may suffer from the issue of data sparsity.